• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name septum_auto

    Log entry time 16:35:42 on April 17, 2012

    Entry number 375913

    keyword=Wrong septum current, please check!

    Left septum set current 100A is not correct, should be 314.898A! please check!
    Left septum read back current 99.3059A is not correct, should be 100A! please check!
    Right septum set current 100A is not correct, should be 412.641A! please check!




    septum alarm using halog_post [halog_post 1.1]

    A copy of this log entry has been emailed to: jixie@jlab.org